Suppose that a committee of 3 people is chosen from 10 people, including Josh. (a) How many total committees are there? (b) How many committees include Josh? (c) Does it seem likely Josh will be chosen for the committee if it is done randomly? Explain
Answer #1

a) Total commitee =  10C3 = 120

b) No. of committees including josh = 1* 9C2 =36

C) P ( josh will be in the committee) = 36/ 120 = 0.3

YES, IT SEEMS LIKELY JOSH WILL BE CHOSEN FOR THE COMMITTE